a golf tournament gave $270000 in prizes. the winner recieved 1/2. the runner up got 1/4. the remainder was divided equally between the third-, fourth, and fifthe -place finishers. how much did the fifth place finisher recieve?
Answer by mathsmiles(68)

A real life problem I wish I had!!

Total prize is $270,000.

Winner receives 1/2 so 270,000/2 = 135,000
That leaves another 135,000 remaining

Runner up (2nd place) receives 1/4 so 270,000/4 = 67500
That leaves 270,000 - 135,000 - 67,500 = 67500

The next 3 places receive this amount divided equally among them.
67500/3 = $22,500

Going back to the original problem, how much did the 5th place finisher receive?
$22,500.

I'd be happy with 5th place! Would you settle for 5th? ;-)